OpenStudy (anonymous):

how do i solve this Y = -16x^2 + 340x + 4500

OpenStudy (anonymous):

Solve it for what? Do you mean you have to solve\[-16x^2+340x+4500=0\]?

OpenStudy (anonymous):

Or do you have a particular x-value and you need to find y?

OpenStudy (anonymous):

Or do you have a y-value and you need to find x?

OpenStudy (anonymous):

5200

OpenStudy (anonymous):

Are you saying y=5200?

OpenStudy (anonymous):

here is the whole problem

OpenStudy (anonymous):

Well, it seems reasonable that y=5200 and you need to find the x-values that make it so.

OpenStudy (anonymous):

2) Snoopy is in his plane, 4500 feet above the ground. He fires a projectile straight up. The initial velocity of the projectile is 340 feet per second. (After the projectile is fired, it has no acceleration, so its height can be modeled by a “falling body” model.) The Red Baron is maintaining a constant altitude of 5200 feet.

OpenStudy (anonymous):

a. Write an equation describing the upward motion of the projectile. (Assume there is no air resistance. Let y represent the altitude of the projectile in feet, and let t represent the time after firing in seconds.) • Y = -16x^2 + 340x + 4500

OpenStudy (anonymous):

c. When, on the way up, might the projectile intersect the Red Baron

OpenStudy (anonymous):

d. If the projectile misses the Red Baron on the way up, when might it intersect him on the way down?

OpenStudy (anonymous):

Is the equation for y an equation you found or were you given that?

OpenStudy (anonymous):

i found the equation

OpenStudy (anonymous):

Well, the problem will depend on the coordinate system you use. If you 'attach' a coordinate system to Snoopy, then when he fires the projectile straight up, it will look like (in your coordinate system) the object is being fired straight up the y-axis. I'm assuming the Red Barron is directly above Snoopy and travelling with same velocity. Does this agree with you so far?

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

Your equation is correct...so now to finish...

OpenStudy (anonymous):

y will be 5200, since your coordinate system is based on the ground, though moving along with Snoopy, so you have to solve for t in:\[5200=-16t^2+340t+4500\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

There will be two times: the first is the initial impact, the second when it comes back down. It has the same momentum in the x-direction as Snoopy and the Red Barron.

OpenStudy (anonymous):

So there are you answers for c and d.

OpenStudy (anonymous):

okay so how do i find out the two times it crosses the red barons path

OpenStudy (anonymous):

\[t=\frac{5}{8}(17\pm \sqrt{177})\]

OpenStudy (anonymous):

You solve the quadratic equation.

OpenStudy (anonymous):

\[5200=-16t^2+340t+4500 \rightarrow -16t^2+340t-700=0\]

OpenStudy (anonymous):

\[t = \frac{-340 \pm \sqrt{340^2-4(-16)(-700)}}{2(-16)}=\frac{-340 \pm \sqrt{70800}}{-32}s\]

OpenStudy (anonymous):

using the quadratic formula (or you can complete the square).

OpenStudy (anonymous):

\[\sqrt{70800}=20\sqrt{177}\]

OpenStudy (anonymous):

Then just simplify to get the answers above. The lower time will be the first strike (answer c), the larger time will be the second strike (answer d).

OpenStudy (anonymous):

ok